OpenStudy (anonymous):

one number is five times another number. a third number is twelve more than the larger of the first two. the sum of the numbers 89. find the numbers. some one please answer this!!!!!!

OpenStudy (mattfeury):

You don't need to post this multiple times. Just once please, and someone will help here.

OpenStudy (anonymous):

someone please answer it then

OpenStudy (anonymous):

lucas, are you figuring it out??/

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

good lucas, make it faster i need this stacked :) :P

OpenStudy (anonymous):

5A= B C = 12B A+B+C= 89 A+5A+60A= 89 66A=89 Answers: A = 1.35 B = 6.75 C = 81

OpenStudy (anonymous):

gw 2011, a lil' faster!!!!!

OpenStudy (gw2011):

Let x=the first number Then 5x=the second number And 5x+12=the third number Now adding the above you get: x+5x+5x+12=89 11x+12=89 11x=77 x=7 This is the first number 5x=(5)(7)=35 This is the second number 5x+12=35+12+47 This is the third number

OpenStudy (gw2011):

The third number should be written as 5x+12=35+12=47. The plus 47 should be equals 47.

OpenStudy (anonymous):

gw2011 is right, I read the problem wrong and calculated the third number being 12 times rather than 12 more.

OpenStudy (anonymous):

kk thanks lucas and gw2011 love you lots